One leg of a right triangle is one half as long as the other leg. The hypotension is 8 meters long. Find the exact length of each leg of the triangle.
Answer by jorel555(1290)

Let x = the short leg. The long leg is 2x. And our right triangle is such that:
x^2+(2x)^2=8^2
x^2+4x^2=64
5x^2=64
x=sqrt (64/5)= approximately 3.58 meters for the short leg, 7.16 meters for the long leg.......